OpenStudy (anonymous):

what is the radius of the circle if its diameter is: 1) 46 cm 2) 16.7 m 3) 35 dm 4) 19.5 ft 5) (2x+6) cm

OpenStudy (anonymous):

\[d = 2r\]I know you can multiply and/or divide by 2.

OpenStudy (anonymous):

how about # 5?...

OpenStudy (anonymous):

x+3

OpenStudy (anonymous):

i need answer...

OpenStudy (anonymous):

tnx.. how about \[4\sqrt{5} m\]

OpenStudy (anonymous):

\(2 \sqrt{5} \)

OpenStudy (anonymous):

great answer tnx...

OpenStudy (anonymous):

a chord of a cicle is 24 in. find the radius of the cicle if chord's distance from the center is: 1) 5 in 2) 9 in 3) 4 in

OpenStudy (anonymous):



OpenStudy (anonymous):

Let's take a look at the above drawing. We can see that \[r^2 = \left( C \over 2 \right)^2 + d^2\]

OpenStudy (anonymous):

what is the answer ?....